A certain airplane flies 250 mph in still air. on a windy day, this plane flew 16 miles into the wind in the same amount of time that it flew 54 miles with a tailwind. Find the speed of the wind.
Answer by TimothyLamb(4379)

s = 250
---
s = d / t
t = d / s
---
T = 16/(s - w)
T = 54/(s + w)
---
T = 16/(250 - w)
T = 54/(250 + w)
---
16/(250 - w) = 54/(250 + w)
16(250 + w)/(250 - w)(250 + w) - 54(250 - w)/(250 - w)(250 + w) = 0
16(250 + w) - 54(250 - w) = 0(250 - w)(250 + w)
16(250 + w) - 54(250 - w) = 0
4000 + 16w - 13500 + 54w = 0
70w = 9500
---
w = 135.71429
---
wind speed is 135.7 mph !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php